OpenStudy (anonymous):

Evaluate f'(2) if f(x)=...

OpenStudy (anonymous):

\[\tan^{-1} (\frac{ 1 }{ x })\]

OpenStudy (anonymous):

I got \[-\frac{ 1 }{ 5 }\]

OpenStudy (zarkon):

good

OpenStudy (anonymous):

:D Thanks.

OpenStudy (anonymous):

I think that right because f'x = -1/(x^2 +1) =>> -1/ (2^2) + 1 = -1/5